Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Yi Qiang
Log entry time 00:39:21 on December 07, 2008
Entry number 252352
This entry is a followup to: 252128
Followups:
keyword=Experiment Summary for Dec 5, Owl - Dec 6, Swing
Summary for Dec 5 owl - Dec 6 swing
Smooth data taking.
Beam in Use
shift |
ABU |
BANU |
BNA |
2008/12/05 owl | 5.87 |
1.55 | 0.59 |
2008/12/05 day | 3.55 |
3.27 | 1.17 |
2008/12/05 swing | 6.78 |
0.15 | 1.08 |
2008/12/06 owl | 7.23 |
0.12 | 0.67 |
2008/12/06 day | 6.50 |
0.41 | 1.10 |
2008/12/06 swing | 7.26 |
0.12 | 0.63 |
total | 37.19 |
5.62 | 5.24 |
Main events/issues:
- Access to temperatarily fix TS problem: 00:00 - 01:00, 2008/12/05.
- Production of pi-: 01:00 - 08:30, 2008/12/05.
- Access to change polarity/Fix TS problem: 08:30 - 12:00, 2008/12/05.
- Production of pi+ on Transverse pumping afterwards.
Plans for the next few days:
- Production of pi+ on Transverse pumping.
- Do moller measurement on Monday (Dec,08) afternoon.